solve the following question step bt step in detail.

∫ x²e³ˣ dx
1
Expert's answer
2020-10-13T18:56:54-0400

Use integration by parts:

"\intop x^2e^{3x}dx=\intop x^2d\frac{e^{3x}}{3}="


"=\frac{x^{2}e^{3x}}{3}-\intop\frac{e^{3x}}{3}dx^2=\frac{x^{2}e^{3x}}{3}-\intop\frac{2xe^{3x}}{3}dx="


"=\frac{x^{2}e^{3x}}{3}-\frac{2}{3}\intop\ xe^{3x}dx=\frac{x^{2}e^{3x}}{3}-\frac{2}{3}\intop xd\frac{e^{3x}}{3}="


"=\frac{x^{2}e^{3x}}{3}-\frac{2}{3}(\frac{xe^{3x}}{3}-\intop\frac{e^{3x}}{3}dx)="


"=\frac{x^{2}e^{3x}}{3}-\frac{2xe^{3x}}{9}+\frac{2}{3}\cdot\frac{e^{3x}}{9}+C=\frac{(9x^{2}-6x+2)e^{3x}}{27}+C"